# VCR: Visual Caption Restoration

### v2

@@ 4.4 Human Evaluation
We recruited 7 volunteers to perform human evaluation on a subset of the samples of our datasets. Two out of the seven evaluators are native English speakers, while five are native Chinese speakers who are also fluent in English171717The TOEFL scores of the non-native English-speaking participants range from 102/120 to 112/120.. All volunteers have earned postgraduate degrees majoring in one of the following fields: biology, statistics, computer science, and economics. The evaluations were conducted on a voluntary basis and participants received no rewards.
We gave the volunteer the following instructions: (1) We ask the volunteers to focus on the puzzles. Each example in the hard collection may require 30 seconds to 2 minutes of focused attention; and (2) we ask the volunteers to utilize the context rather than directly brute-force the puzzle.
Every sample is solved by at least 3 volunteers. In English, we release the exact match score in 2 splits: all errors counted (All), and only count errors not related to date and person names (Filtered).
| EN Easy (N = 169) | EN Hard (N = 169) | ZH Easy (N = 188) | ZH Hard (N = 188) | |||||
@@
Refer to Table A to compare all models with human evaluation results using the same test cases.
@@ 5 Related Work
Masked modeling.
Masked language modeling (MLM) introduced by BERT [13] and its autoregressive counterparts by GPT [8] have been the foundations of pre-training modern natural language processing (NLP) models. Both are trained to predict a masked-out proportion based on the context. Similarly, inspired by BERT and GPT, iGPT [9] uses the input sequences of the pixels to predict the next unknown pixels. Masked Autoencoders (MAE) [19] introduce Masked Image Modeling (MIM) by masking out random patches from input images and reconstructing the missing parts in the images in pixel space. The goal is to encourage the model to learn the representations of the image by understanding the context and structure of the visible patches, which helps to reconstruct the masked images accurately.
@@
Several datasets have been proposed for visual question answering VQA [3, 77, 17, 47]. FVQA [65] and OK-VQA[44] are datasets about knowledge-based visual question answering and contains questions that necessitate the usage of external knowledge resources. CLEVR [30] is a synthetic VQA dataset that mainly focuses on visual reasoning abilities. Recognizing the need to develop VQA models that can understand text, Text-VQA [56, 6, 48, 68] aims to read and reason about texts embedded within images in the context of image-question answering. Several datasets [56, 6, 48] have been developed for the Text-VQA task, such as the TextVQA dataset [56] and the ST-VQA dataset [6] on natural images, the OCR-VQA dataset [47] on book or movie covers, the InfographicVQA [45] dataset on infographics, and the DocVQA dataset [46] on document images.
@@
Vision-language models are designed for tasks that involve understanding and generating content from images and text [58, 40, 31, 32]. For example, models have been developed to combine Llama3 with advanced vision-language processing capabilities to handle complex multimodal tasks [75, 71, 24, 74, 67, 14]. Qwen-VL [4] enhances visual-linguistic representations for more accurate contextual interpretations, while OpenGVLab-InternVL-Chat [11, 10] merges the InternVL framework with interactive chat capabilities. These studies typically employ a multimodal encoder [53, 76, 70] to process multimodal data, which is then mapped to the same input space of the language model. General-purpose models such as the GPT-4 series models [51, 50], the Claude series models [2], the Gemini series models [59] and the Reka series models [60] have also been adapted for vision-language tasks, demonstrating strong performance in multimodal tasks. Finally, DocLLM [64] specializes in document understanding by integrating visual and textual data to enhance the interpretation and generation of document-related content. These models collectively represent significant advancements in vision-language integration, contributing unique capabilities and enhancements to the understanding and generation of multimodal information.
@@
OCR [49] and its subproblems [21, 57, 54, 15] have been well-studied in the literature in the constrained setting. However, classical OCR methods often cannot perform well on images captured in the wild in an unconstrained setting. Many new methods have been developed for advancing scene-text recognition on camera-captured images [5, 18, 26, 28, 66, 55, 81, 33]. In addition to the detection and recognition of OCR tasks, visual question answering has emerged as an important downstream task in the OCR literature. With the development of Text-VQA, new methods for improving the reading abilities in VQA utilizing OCR have been proposed. For example, LoRRA [56] extends a VQA model Pythia [29] with an OCR module to better handle Text-VQA tasks. TAP [72] incorporates scene texts that are generated from OCR engines during pretraining to further improve Text-VQA capabilities.
Scene Text Detection (STD).
Identifying and interpreting text in natural images serves as a foundational step towards developing a comprehensive VQA system. This capability allows for the integration of textual information, enhancing overall scene comprehension. Existing approaches mainly consist of two steps: text detection and recognition. Several methods based on Fully Convolutional Neural Networks have been proposed for scene text detection [36, 37, 81, 20, 80]. In terms of text recognition, [27] proposed to directly recognize texts from entire input images using a 90k-class convolutional neural network with each class corresponding to an English word [37]. In addition, several methods [73, 78, 42] have been proposed to tackle the issue of irregular text in scene text detection.
@@ 6 Conclusion
In this work, we introduced the Visual Caption Restoration (VCR) task, a novel vision-language challenge aimed at promoting the integration of visual and textual modalities, including text embedded in both natural language tokens and image formats and highly obscured text embedded in the image. We developed a specialized pipeline to create a dataset tailored to this task, utilizing correlated image-text pairs. This task stands out from existing methods by requiring a more profound integration of visual cues and partially obscured text, highlighting its uniqueness and importance in the field.
@@
By introducing the VCR task and its specialized dataset, we aim to advance research in vision-language interaction. The unique challenges of VCR seek to improve model development and training, extending the limits of multimodal AI. We invite the community to utilize our dataset and develop innovative strategies to boost the performance of vision-language models.
Acknowledgments and Disclosure of Funding
We acknowledge the contribution of Mila for supporting the evaluation of various models on our datasets through providing GPU access. None of the authors received third-party funding during the last 36 months prior to this submission.
@@ Appendix A Additional evaluation results on first 100 and 500 test cases
English Easy Closed-source models
Claude 3 Opus - -20 -13.06
@@
InternVL-V1.5 25.5B -59.76 -33.66
@@
Hard Closed-source models
Claude 3 Opus - -17 -13.31
@@
InternVL-V1.5 25.5B -5.33 -10.32
@@
Chinese Easy Closed-source models
Claude 3 Opus 0 2.2
@@
Open-source models
CogVLM2-Chinese 19B 0.53 1.52
@@
InternVL-V1.5 25.5B 1.06 10.59
@@
Hard Closed-source models
Claude 3 Opus - 0.53 1.45
@@
Open-source models
CogVLM2-Chinese 19B 0 -3.05
@@
InternVL-V1.5 25.5B 0 3.03
@@
English Easy Closed-source models
Claude 3 Opus - -15 -10.74
@@
InternVL-V1.5 25.5B -59.13 -34.82
@@
Hard Closed-source models
Claude 3 Opus - -12.2 -12.48
@@
InternVL-V1.5 25.5B -4.6 -9.26
@@
Chinese Easy Closed-source models
Claude 3 Opus - -0.1 1.49
@@
Open-source models
CogVLM2-Chinese 19B 2.2 3.92
@@
InternVL-V1.5 25.5B -0.7 5.15
@@
Hard Closed-source models
Claude 3 Opus - 0.2 1.13
@@
Open-source models
CogVLM2-Chinese 19B -1.1 -3.04
@@
InternVL-V1.5 25.5B 0 1.79
@@
We show the table of evaluation results on first 100 and 500 test cases for better comparison with human evaluation results and closed-source models correspondingly.

### v3

@@ 4.5 Human Evaluation
We recruited 7 volunteers to perform human evaluation on a subset of the samples from our datasets. Two out of the seven evaluators are native English speakers, while five are native Chinese speakers who are also fluent in English444The TOEFL scores of the non-native English-speaking participants range from 102/120 to 112/120.. All volunteers have earned postgraduate degrees, majoring in one of the following fields: biology, statistics, computer science, and economics. The evaluations were conducted on a voluntary basis, and participants received no rewards.
We gave the volunteers the following instructions: (1) We asked the volunteers to focus on the puzzles. Each example in the hard collection may require 30 seconds to 2 minutes of focused attention, and (2) we asked the volunteers to utilize the context rather than directly brute-force the puzzle.
Every sample is solved by at least 3 volunteers. In English, we release the exact match score in 2 splits: all errors counted (All), and only counting errors not related to dates and person names (Filtered).
The human evaluation results are shown in Table 3. Although current SOTA models suffer from the challenge, fluent speakers can easily achieve more than 90 percent accuracy across difficulties. Please refer to Table A to compare all models with human evaluation results using the same test cases.
| (N = 169) | (N = 169) | (N = 188) | (N = 188) | |||||
@@ 5 Related Work
Several datasets have been proposed for visual question answering VQA Antol et al. (2015); Zhang et al. (2016); Goyal et al. (2017); Mishra et al. (2019b). FVQA Wang et al. (2018) and OK-VQAMarino et al. (2019) are datasets about knowledge-based visual question answering and contains questions that necessitate the usage of external knowledge resources. CLEVR Johnson et al. (2017) is a synthetic VQA dataset that mainly focuses on visual reasoning abilities. Recognizing the need to develop VQA models that can understand text, Text-VQA Singh et al. (2019); Biten et al. (2019); Mishra et al. (2019a); Wang et al. (2020) aims to read and reason about texts embedded within images in the context of image-question answering. Several datasets Singh et al. (2019); Biten et al. (2019); Mishra et al. (2019a) have been developed for the Text-VQA task, such as the TextVQA dataset Singh et al. (2019) and the ST-VQA dataset Biten et al. (2019) on natural images, the OCR-VQA dataset Mishra et al. (2019b) on book or movie covers, the InfographicVQA Mathew et al. (2022) dataset on infographics, and the DocVQA dataset Mathew et al. (2021) on document images.
@@
Vision-language models are designed for tasks that involve understanding and generating content from images and text Sun et al. (2023); Liu et al. (2023b); Laurençon et al. (2023; 2024). For example, models have been developed to combine Llama3 with advanced vision-language processing capabilities to handle complex multimodal tasks Yu et al. (2024); Xu et al. (2024); Hu et al. (2023); Yu et al. (2023); Wang et al. (2023b); Dong et al. (2024a). Qwen-VL Bai et al. (2023) enhances visual-linguistic representations for more accurate contextual interpretations, while OpenGVLab-InternVL-Chat Chen et al. (2023; 2024b) merges the InternVL framework with interactive chat capabilities. These studies typically employ a multimodal encoder (Radford et al., 2021; Zhai et al., 2023; Wu et al., 2022) to process multimodal data, which is then mapped to the same input space of the language model. General-purpose models such as the GPT-4 series models (Ouyang et al., 2022; OpenAI et al., 2023), the Claude series models (Anthropic, 2024), the Gemini series models (Team et al., 2024a) and the Reka series models (Team et al., 2024b) have also been adapted for vision-language tasks, demonstrating strong performance in multimodal tasks. Finally, DocLLM Wang et al. (2023a) specializes in document understanding by integrating visual and textual data to enhance the interpretation and generation of document-related content. These models collectively represent significant advancements in vision-language integration, contributing unique capabilities and enhancements to the understanding and generation of multimodal information.
@@
OCR Nagy (2000) and its subproblems Howe (2013); Smith (1995); Shafait et al. (2008); Frinken et al. (2011) have been well-studied in the literature in the constrained setting. However, classical OCR methods often cannot perform well on images captured in the wild in an unconstrained setting. Many new methods have been developed for advancing scene-text recognition on camera-captured images Bissacco et al. (2013); Gupta et al. (2016); Huang et al. (2014); Jaderberg et al. (2014); Wang et al. (2012); Shi et al. (2017); Zhou et al. (2017); Lee & Osindero (2016). In addition to the detection and recognition of OCR tasks, visual question answering has emerged as an important downstream task in the OCR literature. With the development of Text-VQA, new methods for improving the reading abilities in VQA utilizing OCR have been proposed. For example, LoRRA Singh et al. (2019) extends a VQA model Pythia Jiang et al. (2018) with an OCR module to better handle Text-VQA tasks. TAP Yang et al. (2021) incorporates scene texts that are generated from OCR engines during pretraining to further improve Text-VQA capabilities.
@@ 6 Conclusion
In this work, we introduced the VCR task, a novel vision-language challenge aimed at promoting the integration of visual and textual modalities, including text embedded in both natural language tokens and image formats and highly obscured text embedded in the image. We developed a specialized pipeline to create a dataset tailored to this task, utilizing correlated image-text pairs. This task stands out from existing methods by requiring a more profound integration of visual cues and partially obscured text, highlighting its uniqueness and importance in the field.
@@
By introducing the VCR task and its specialized dataset, we aim to advance research in vision-language interaction. The unique challenges of VCR seek to improve model development and training, extending the limits of multimodal AI. VCR provides a controllable testbed for fine-grained analysis of model behavior across languages, difficulty levels, image inclusion, and fine-tuning stages. We invite the community to utilize our dataset and develop innovative strategies to boost the performance of vision-language models.
@@ Appendix A Additional evaluation results on first 100 and 500 test cases
English Easy Closed Claude 3 Opus - -20 -13.06
@@
InternLM-XComposer2-VL-4K 7B 0.59 0.06
InternLM-XComposer2.5-VL 7B 16.57 13.45
InternVL-V2 40B -0.59 -0.84
Llama-3.2-Vision 11B 11.24 7.97
Llama-3.2-Vision 90B 8.28 4.64
@@
Hard Closed Claude 3 Opus - -17 -13.31
@@
InternLM-XComposer2-VL-4K 7B -0.59 0.84
InternLM-XComposer2.5-VL 7B -1.18 -2.48
InternVL-V2 40B -4.14 -5.76
Llama-3.2-Vision 11B 2.96 6.70
Llama-3.2-Vision 90B -2.37 -3.06
@@
Chinese Easy Closed Claude 3 Opus - 0 2.2
@@
Open CogVLM2-Chinese 19B 0.53 1.52
@@
InternLM-XComposer2-VL-4K 7B 0.00 -1.08
InternLM-XComposer2.5-VL 7B -1.60 -4.18
InternVL-V2 40B 6.91 7.72
@@
Hard Closed Claude 3 Opus - 0.53 1.45
@@
Open CogVLM2-Chinese 19B 0 -3.05
@@
InternLM-XComposer2-VL-4K 7B 0.00 -0.39
InternLM-XComposer2.5-VL 7B 0.00 0.32
InternVL-V2 40B -0.53 -1.32
@@
English Easy Closed Claude 3 Opus - -15 -10.74
@@
InternLM-XComposer2-VL-4K 7B 1.99 4.03
InternLM-XComposer2.5-VL 7B 16.65 12.28
InternVL-V2 40B -2.24 -1.71
Llama-3.2-Vision 11B 12.80 9.08
Llama-3.2-Vision 90B 9.32 5.29
@@
Hard Closed Claude 3 Opus - -12.2 -12.48
@@
InternLM-XComposer2-VL-4K 7B -0.12 0.37
InternLM-XComposer2.5-VL 7B -0.50 -1.25
InternVL-V2 40B -4.35 -6.02
Llama-3.2-Vision 11B 6.58 9.64
Llama-3.2-Vision 90B 2.98 2.43
@@
Chinese Easy Closed Claude 3 Opus - -0.1 1.49
@@
Open CogVLM2-Chinese 19B 2.2 3.92
@@
InternLM-XComposer2-VL-4K 7B 0.10 -1.74
InternLM-XComposer2.5-VL 7B -0.10 -2.23
InternVL-V2 40B 6.09 10.05
@@
Hard Closed Claude 3 Opus - 0.2 1.13
@@
Open CogVLM2-Chinese 19B -1.1 -3.04
@@
InternLM-XComposer2-VL-4K 7B 0.00 0.25
InternLM-XComposer2.5-VL 7B 0.00 0.02
InternVL-V2 40B -0.50 -1.50
@@
Appendix B More relationship between VCR and other benchmarks
The heatmap shown in Figure 6 provides a detailed view of the pairwise correlation between 23 different benchmarks used to evaluate 38 VLMs. The scores of these models on each benchmark were utilized to compute this correlation matrix. The color intensity and numerical values represent the degree of correlation, ranging from -1 (perfect negative correlation) to 1 (perfect positive correlation), with warmer colors indicating higher positive correlations and cooler colors indicating weaker or negative correlations.
We observe that is markedly different from other benchmarks in the evaluation set. It demonstrates minimal and even sometimes negative correlation with other benchmarks. This suggests that the skill set required for is largely unrelated to those tested by other popular tasks emphasizing more straightforward image-to-text associations or OCR capabilities. Specifically, challenges models with tasks that involve high-level reasoning and minimal reliance on pixel-level information, focusing instead on understanding context, commonsense reasoning, and visual narrative interpretation. These features are less critical in other benchmarks, which explains the weak correlation across tasks. exhibits a slightly stronger correlation with a few other benchmarks but remains moderately independent of most others. Like , also evaluates visual commonsense reasoning, but with less stringent requirements, offering models more cues and simpler connections between visual elements and textual understanding. This leads to moderate overlap with benchmarks like TextVQA, which similarly focus on text understanding in a visual context, but still emphasizes a higher level of interpretative reasoning than standard text-based vision benchmarks.
Appendix C Information of models evaluated
We provide the specifications of the evaluated models and a link to each model’s weights in Table 6.
Appendix D Potential QA
What could be the possible reason that CogVLM performs well in VCR-wiki series benchmarks?
Many models we tested (DocOwl-1.5, Monkey, MiniCPM-V2.5, InternLM series, InternVL series) follow a similar inference pipeline to adapt to high-resolution application scenarios:
1.
An algorithm divides the input image into segments.
2.
Each segment is encoded into tokens using a CILP-based image encoder.
3.
A filtering mechanism (algorithm/resampler/abstractor) processes the visual tokens.
4.
The filtered tokens are concatenated with language tokens and input to the LLM.
If, in step 3, pixel-level hints embedded in text within the image (TEL) are disregarded, the model cannot correctly answer the question. Consequently, some of these models may perform better on benchmarks emphasizing global features but struggle on the VCR-wiki series benchmarks, particularly in the hard partitions. For example, while InternVL2-40B performs best on , it does not perform well on . As noted in the paper, the easy partition of the benchmark primarily verifies that the VCR task is feasible for the models. In contrast, the hard partition explores the boundaries of VCR capability for both models and human test-takers (who require more time and focus to solve the puzzles in the hard partition).
The CogVLM2 and Cambrian-1 series, by contrast, do not include step 3 in their inference pipelines. Instead, their image encoders operate at mid-to-high resolutions (1K level), and they resize the input image to match the supported resolution rather than dividing it into segments. The image encoder resolution for CogVLM2 is , while Cambrian-1 employs four image encoders, the largest supporting a resolution of . This approach may encounter challenges with extremely shaped input images (e.g., ), but for VCR-wiki, where images are mostly near-square (on average for / and for /), high-resolution support is not necessary. For instance, InternLM-XComposer2-VL outperforms InternLM-XComposer2-VL-4KHD on this benchmark.
What could be the potential way to improve models’ capability on VCR?
To suggest potential avenues for improving VLM performance on VCR, we propose the following:
1.
Include VCR in VLM pretraining: Just as OCR parsing tasks are often included in pretraining to improve OCR performance, researchers could consider incorporating VCR tasks during pretraining. We have released vcr_transform.py242424https://github.com/tianyu-z/VCR/blob/main/src/build_dataset/vcr_transform.py to facilitate this process, making it as straightforward as data augmentation.
2.
Architectural Exploration: CogVLM2 is the best-performing model on average across the four partitions, and we believe this is largely due to its vision expert architecture. We contacted the CogVLM2 team and learned that GLM-4 and CogVLM2 share the same training data, yet there is a significant performance gap between them on the VCR benchmarks.
3.
Chain-of-Thought (CoT) Methods: Researchers could explore multi-modality pipelines based on CoT techniques to improve existing VLMs on VCR tasks Chen et al. (2024a); Zhang et al. (2023). Although a model might not initially focus on the correct visual area (e.g., pixel-level hints in the TEI), CoT-based techniques could help refine its focus over successive rounds.
How is the Transferability of VCR-wiki Finetuning?
In Table 7, we show the transferability of VCR-wiki by finetuning multiple models on different finetuning datasets’ training sets and testing their performance on a series of benchmarks.
The analysis of our experimental results highlights the strong transferability of the proposed VCR-wiki dataset across various benchmarks. Notably, models fine-tuned on VCR-wiki demonstrate significant performance improvements not only within the VCR-wiki benchmarks themselves, but also across different language settings. For example, fine-tuning CogVLM2 on leads to a substantial increase in performance on the benchmark, elevating the score from 9.15 to 42.55. Similarly, fine-tuning the Chinese version of CogVLM2 on enhances its performance on both the and benchmarks, with scores rising from 79.9 to 87.57 and from 25.13 to 44.97, respectively. These enhancements indicate that the VCR-wiki dataset facilitates the learning of effective transferable features even when the fine-tuning and evaluation involve different languages.
Additionally, the consistent achievement of the highest scores within each model’s finetuning variations underscores the robustness of VCR-wiki in improving model performance across diverse evaluation metrics. This evidence collectively demonstrates that VCR-wiki serves as a versatile and powerful resource for enhancing model generalization and performance across multiple tasks and linguistic contexts.
| Base Model | MiniCPM-V2.5 | MiniCPM-V2.5 | MiniCPM-V2.5 | MiniCPM-V2.5 | CogVLM2 | CogVLM2 | CogVLM2 | CogVLM2-Ch. | CogVLM2-Ch. | CogVLM2-Ch. |
| Finetuning_dataset | None | OKVQA-Train | None | OKVQA-Train | None | OKVQA-Train | ||||
| OKVQA | 77.43 | 72.20 | 77.09 | 76.38 | 75.35 | 71.86 | 75.45 | 74.16 | 70.57 | 74.14 |
| 31.81 | 19.53 | 40.96 | 28.40 | 83.25 | 79.29 | 93.27 | 79.90 | 30.18 | 87.57 | |
| 1.41 | 0.00 | 13.86 | 5.33 | 37.98 | 27.22 | 77.44 | 25.13 | 3.55 | 44.97 | |
| 4.10 | 2.12 | 2.66 | 7.44 | 9.15 | 16.49 | 42.55 | 33.24 | 16.49 | 61.69 | |
| 0.09 | 0.53 | 1.06 | 1.53 | 0.08 | 0.00 | 1.60 | 1.34 | 0.00 | 42.11 | |
| MMstar | 50.20 | 51.73 | 50.40 | 50.27 | 50.50 | 51.07 | 50.20 | 52.73 | 50.87 | 54.33 |
| MMBench_DEV_EN | 74.54 | 74.46 | 74.54 | 74.30 | 72.70 | 73.53 | 72.60 | 77.32 | 77.09 | 76.78 |
| MME | 2024.6 | 1996.7 | 1923.65 | 1977.13 | 1869.5 | 1860.56 | 1882.21 | 2040.7 | 1898.42 | 1939.8 |
| MMMU_VAL | 45.89 | 47.78 | 46.11 | 46.56 | 42.60 | 38.67 | 38.67 | 42.44 | 41.56 | 45.00 |
| AI2D_test | 78.04 | 77.85 | 77.62 | 77.91 | 73.40 | 74.97 | 73.61 | 72.64 | 70.98 | 71.31 |
| OCR_BENCH | 71.70 | 71.60 | 71.50 | 71.30 | 75.40 | 72.80 | 79.80 | 77.30 | 75.20 | 79.60 |
| MMVet | 53.12 | 45.78 | 51.10 | 52.66 | 57.80 | 45.00 | 59.31 | 56.38 | 39.77 | 56.88 |
| MathVista_MINI | 54.50 | 53.70 | 52.80 | 52.70 | 38.60 | 38.30 | 35.40 | 37.80 | 37.60 | 40.20 |
| ChartQA | 71.80 | 72.12 | 71.92 | 72.52 | 72.80 | 80.40 | 79.56 | 63.16 | 58.92 | 65.84 |
| OCRVQA | 61.85 | 62.70 | 61.36 | 61.59 | 64.90 | 65.56 | 66.11 | 32.65 | 34.41 | 33.24 |